Solving the formula L=2d+pi(a+r) for a.
Answer by nerdybill(7384)

The idea is to isolate 'a' to one side of the equation -- doesn't matter which side.
.
L=2d+pi(a+r)
Start by subtracting 2d from both sides to get:
L-2d = pi(a+r)
.
Now, divide both sides by pi:
(L-2d)/pi = a+r
.
Finally subtract r from both sides:
((L-2d)/pi) - r = a